OpenStudy (anonymous):

solve 1 1/3 t = 10 will give metal to best answer.

OpenStudy (anonymous):

\[1 \frac{ 1 }{ 3 } t = 10\]

OpenStudy (austinl):

I would first get that fraction as a complex fraction.

OpenStudy (anonymous):

i don't have that

OpenStudy (austinl):

For example, \(\displaystyle 1\frac{2}{5}=\frac{7}{5}\)

OpenStudy (anonymous):

i know. i don't have it.

OpenStudy (anonymous):

do you know the answer or not?

OpenStudy (anonymous):

\[1+\frac{1}{3}=\frac{4}{3},\frac{4}{3}t=10,t=10\left(\frac{3}{4}\right) \]

OpenStudy (anonymous):

thanks but i don't think that is right

OpenStudy (anonymous):

i think it would be \[16 \frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

or \[7 \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

not sure what one

OpenStudy (anonymous):

im pretty sure its 7.5

OpenStudy (anonymous):

or 7 1/2

OpenStudy (anonymous):

fraction or decimal

OpenStudy (anonymous):

\[10\left(\frac{3}{4}\right)=\frac{15}{2} \]